OpenStudy (anonymous):

A particle is moving along a projectile path where the initial height is 160 feet with an initial speed of 144 feet per second. What is the maximum height of the particle? 128 feet 224 feet 272 feet 484 feet

OpenStudy (anonymous):

Assuming Earth's gravity, the formula for the flight of the particle is: s(t) = -16t^2 + vt + s = -16t^2 + 144t + 160. This has a maximum when t = -b/(2a) = -144/(-32) = 9/2. Therefore, the maximum height is s(9/2) = -16(9/2)^2 + 144(9/2) + 160 = 484 ft. I hope this helps!

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

hahahh elvin.. shut up (:

OpenStudy (anonymous):

what i do? :)

OpenStudy (anonymous):

uhh, nothing?

OpenStudy (anonymous):

uhh,nothing?

OpenStudy (anonymous):

MUHAHAHAHAHAHAH!!!!

OpenStudy (anonymous):

yeah i knowwwww hahaha

OpenStudy (anonymous):

lol